A 1.5m wide door (mass 15kg) is hinged at one side so that it can rotate without friction about a vertical axis. It is unlatc1. Find angular speed of door after collision

2. What is the change in kinetic friction?
Answer #1

by angular momentum conservation

Li = Lf

mv*L/2 = ( 1/3*ML2 +m(L/2)2 ) ω

10*10^-3* 400*1.5/2 = ( 1/3*15*1.5^2+10*10^-3*(1.5/2)^2 ) ω

10.10-9 _400-101+ ( < 15x1.52 + 10 109 (5) Result: 3 = 11.2556 w

from here solve for ω

ω≈0.266533 rad/s answer

2)

change in kinetic energy

KE f - KE i

= 1/2*I*ω2 - 1/2*m*v2

=0.5*( 1/3*15*1.5^2+10*10^-3*(1.5/2)^2 )*0.266533^2 - 1/2*10*10^-3*400^2 =

05(ž 15x152 +10-10° (* 0.266533 - 10-10-* 4002 Result: -799.6002010999491246875

change in kinetic energy = -799.6 J answer